A container contains water upto a height of 20cm and there is a point source at the centre of the container. The ceiling of the room is 2m above the water surface. Refractive index of water is 4/3. A ring of radius r is placed on water surface such that its centre lies vertically above the surface. Find the radius of shadow of the ring formed on the ceiling if radius of the ring is 15cm?

Grade:12

1 Answers

Debojyoti Das
6 years ago
r = R + h/√u^2 - 1, u = mewr = 0.15 + 2.2/0.88(approx) Now u canPut the values, u will get the answer I hope u will get the answer